OpenStudy (anonymous):

After a number is increased by 1/3 of its value, the result is 664. What was the original value?? Thank you!! please help!

OpenStudy (anonymous):

Let us say the number is \(x\). Then for it to be increased by \(\frac 13\) its value and have that value equal \(664\) means the following.\[x+\frac{x}{3}=664\]Now, just solve for \(x\).

OpenStudy (anonymous):

would i have to do a square root?

OpenStudy (anonymous):

oh no, i would just have to divide be 2 right?

OpenStudy (anonymous):

you should eventually get 498 by solving that

OpenStudy (anonymous):

oh thank you for your answer

OpenStudy (anonymous):

did you figure it out?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

take x+(1/3)x = 664 factor out the x to get x(1+1/3) = 664 x(4/3) = 664 x= (664*3)/4 x=498

OpenStudy (anonymous):

thank you very much. Sorry for my ignorance